OpenStudy (anonymous):

how much work is performed when you push a couch of mass 44kg a distance of 2.5 m using a force of 135 N?

OpenStudy (anonymous):

What is the definition of work?

OpenStudy (anonymous):

Hint: \[W = F d \cos \theta\]

OpenStudy (anonymous):

I think theyre looking for work in joules

OpenStudy (anonymous):

Yes, the units of work is in Joules, I gave you the formula, above.

OpenStudy (shamim):

If theta=0 W=?

OpenStudy (shamim):

Cos0=?

OpenStudy (anonymous):

If it doesn't mention that the box is pushed up a hill or down a hill etc, that means it's completely parallel to the floor, meaning the angle is 0. (cos(0) = 1)

OpenStudy (anonymous):

wait so what would the answer be..?

OpenStudy (shamim):

Force F=135N Displacement s=2.5m

OpenStudy (shamim):

Work w=135*2.5Joule